# RDD: Robust Feature Detector and Descriptor using Deformable Transformer

### v4

@@ 1 Introduction
Keypoint detection and description are central to numerous 3D computer vision tasks, including structure-from-motion (SfM), visual localization, and simultaneous localization and mapping (SLAM). These tasks depend on reliable keypoints and descriptors to establish accurate tracks, which are crucial for downstream algorithms to interpret spatial relationships—a requirement in fields from robotics to augmented reality [20]. When applied to real-world tasks, these systems face challenges such as significant viewpoint shifts, lighting variations, and scale changes.
@@
Historically, feature detection and description methods relied on rule-based techniques such as SIFT , SURF, and ORB [25, 3, 36]. However, these methods oftenf struggled with the aforementioned challenges. To better address these issues, recent works [6, 13, 44, 49] leverage deep neural networks based on data-driven approaches which often extract more robust and discriminative descriptors than the rule-based ones. Despite the enhanced robustness that learning-based methods have, they often rely on regular convolutions to encode images [6, 44, 13, 33, 49], which ignore the geometric invariance and long-distance awareness critical for robust description under challenging conditions. Subsequently, [28, 48, 26] have partially addressed geometric invariance by estimating the scale and orientation of keypoint descriptors to model affine transformations. ALIKED [50] and ASLFeat [28] advanced this approach by using deformable convolutions capable of modeling any geometric transformations. However, these methods are limited to modeling transformations within local windows; features learned with the localized kernels of convolution operators fall short of learning important visual cues that depend on long-range relationships (e.g. vanishing lines).
@@
In this paper, we focus on feature detection and description in challenging scenarios that haven’t been effectively addressed in prior works – specifically, extracting reliable and discriminative keypoints and descriptors across images with large camera baselines, significant illumination variations, and scale changes. We therefore propose RDD: a novel two-branch architecture for keypoint detection and description that can model both geometric invariance and global context at the same time. In particular, we design two dedicated network branches: a fully convolutional architecture to perform keypoint detection and a transformer-based architecture for extracting descriptors: the objectives of keypoint detection and description have previously been found not to align perfectly [21]. We observe that the convolutional neural network excels in detecting sub-pixel keypoints thanks to its expressiveness, whereas a transformer-based architecture is particularly suited for learning features that captures global context and possess geometric invariance through its self-attention mechanism [45, 7]. However, as self-attention layers attend to all spatial locations in the image, they significantly increase computational overhead and may reduce descriptor discriminability. To address this, we have adopted the deformable attention [52] for keypoint description, allowing our designed network to selectively focus on key locations, which greatly reduces time complexity while maintaining the capability to learn geometric invariance and global context.
@@
The two-branch architecture of RDD allows specialized architectures to learn keypoint detection and description independently. Our experiments have found that such design choice accelerates convergence and leads to better overall performance, as proven in Sec. 4.4. Overall, RDD is designed to be robust, achieving competitive performance on challenging imagery and also capable of semi-dense matching Fig. 1. This enhanced capability improves accuracy and stability in 3D vision applications, such as structure-from-motion (SfM) [4] and relative camera pose estimation under difficult imaging conditions. Our approach outperforms current state-of-the-art methods for keypoint detection and description on standard benchmarks, including MegaDepth-1500 [8], HPatches [1], and Aachen-Day-Night [39], across various tasks. However, to the best of our knowledge, no existing benchmarks accurately capture the aftermentioned challenging scenarios. To further validate the robustness of our method and provide a more comprehensive evaluation, we have collected and derived two additional benchmark datasets. The main contributions of our work are:
@@
  • Dataset contribution: MegaDepth-View, derived from MegaDepth [22], specifically focuses on challenging image pairs with large viewpoint and scale variations. Air-to-Ground dataset provides diverse aerial and ground cross view imagery for training and evaluation, enhancing robustness and performance and providing a new evaluation setting.
Figure 2: An overview of our network architecture. Descriptor Branch $\mathcal{F}_{D}$ and Keypoint Branch $\mathcal{F}_{K}$ process an input image $I\in\mathbb{R}^{H\times W\times 3}$ independently. Descriptor Branch: 4 layers of multiscale feature maps $\{\mathbf{x}_{res}^{l}\}_{l=1}^{L}$ are extracted by passing $I$ through ResNet-50 [15] $\mathcal{F}_{res}$. Additional feature map at scale of $1/64$ is added by applying a simple CNN on the last feature map and then they are feed to a transformer encoder $\mathcal{F}_{e}$ with positional embeddings [45]. We up-sample all feature maps output by $\mathcal{F}_{e}$ to size $H/k\times W/k$ where $k=4$ is the patch size. Feature maps are then summed together to generate the dense descriptor map $D$. A classification head $\mathcal{F}_{m}$ is applied to the descriptor map to estimate a matchability map $M$. Keypoint Branch: $I$ passes through a lightweight CNN with residual connection [15] $\mathcal{F}_{cnn}$ to capture multi-scale features $\{\mathbf{x}_{cnn}^{l}\}_{l=1}^{L}$. Features are then up-sampled to size $H\times W$ and concatenated to generate a feature map of $H\times W\times 64$. A score map $S$ is estimated by a classification head $\mathcal{F}_{s}$. Final sub-pixel keypoints are detected using DKD [49].
@@ 2.1 Geometric Invariant Descriptors
In geometric invariance modeling, previous methods have primarily focused on two key aspects: scale and orientation. In traditional hand-crafted methods, SIFT [25] estimates a keypoint’s scale and orientation by analyzing histograms of image gradients and subsequently extracts and normalizes image patches around the keypoint to construct scale- and orientation-invariant descriptors. Similarly, ORB [36] computes orientation based on the keypoint’s center of mass, followed by rotation of the image patches to ensure orientation invariance.
Learning-based methods approach geometric invariance through two primary strategies. The majority of these methods [6, 44, 8, 34, 13, 33] relies heavily on data augmentation techniques to achieve scale and orientation invariance. On the other hand, some learning-based approaches explicitly model scale and orientation. For instance, LIFT [48] mimics SIFT [25] by detecting keypoints, estimating their orientation, and extracting descriptors with different neural networks. Other methods, such as AffNet [30] and LF-Net [31], predict affine transformation parameters and apply these transformations via Spatial Transformer Networks [18] to produce affine-invariant descriptors. These methods generally assume a predefined geometric transformation, typically affine transformations. More recent advancements, such as ASLFeat [28] and ALIKED [50], employ deformable convolutions to directly extract geometric-invariant features, offering a more flexible and adaptive approach to handling geometric variations.
@@ 2.2 Decoupling Keypoint and Descriptor Learning
In traditional handcrafted methods for feature detection and description, such as SIFT [25] and ORB [36], the detection and description processes are performed independently. This decoupled approach was also common in early learning-based methods [2, 29], where keypoint detection and description were treated as separate tasks. However, later works [8, 6, 13, 34] proposed an end-to-end framework, jointly optimizing keypoint detection and descriptor extraction. This joint learning approach ensures that the detected keypoints and their descriptors complement each other. Despite the advantages,  [21] highlighted a potential drawback of joint training, observing that poorly learned descriptors could negatively impact keypoint detection quality. Subsequently, DeDoDe[10] demonstrated that decoupling keypoint detection and description into two independent networks can significantly improve performance. Similarly, XFeat [33] showed that using task-specific networks provides a better balance between efficiency and accuracy. Inspired by these recent insights, RDD consists of two separate branches for independent keypoint detection and description. We also formulate a training strategy that optimize the descriptor branch first and then the keypoint branch. This design enables both sparse and dense matching and ensures optimized performance across various scenarios.
@@ 2.3 Efficient Attention Mechanisms
The quadratic complexity of multi-head attention [45] leads to long training schedules and limited spatial resolution for tractability. For 2D images, LR-Net [16] proposed a local relation layer, computing attention in a fixed local window. Linear Transformer [19] proposes to reduce the computation complexity of vanilla attention to linear complexity by substituting the exponential kernel with an alternative kernel function. Deformable attention [52] further allows the attention window to depend on the query features. With a linear projection predicting a deformable set of displacements from the center and their weight, each pixel in the feature map may attend to a fixed number of pixels of arbitrary distance. Deformable attention [52] is well-suited for our objectives, enabling efficient learning of global context, making it a fitting choice for our approach.
@@ 3 Methods
Our method extracts descriptors and keypoints separately using a descriptor branch and a keypoint branch, as shown in Fig. 2. For an input image $I\in\mathbb{R}^{H\times W\times 3}$, the descriptor branch first extracts multi-scale feature maps $\{x_{res}^{l}\}_{l=1}^{L}$ from $I$ using ResNet-50 [15], denoted as $\mathcal{F}_{res}$, and then estimates a descriptor map $D\in\mathbb{R}^{H/k\times W/k}$ through a deformable transformer encoder [52], denoted as $\mathcal{F}_{e}$, where $k=4$ is the patch size. The matchability map $M\in\mathbb{R}^{H/k\times W/k}$ is then estimated from $D$ with a classification head $\mathcal{F}_{m}$. The keypoint branch $\mathcal{F}_{K}$ estimates a scoremap $S\in\mathbb{R}^{H\times W}$ by passing image $I$ through a lightweight CNN with residual connection [15] $\mathcal{F}_{cnn}$ and a classification head $\mathcal{F}_{s}$, then sub-pixel keypoints $p_{s}=[x,y]^{T}$ are detected with differentiable keypoint detection (DKD) [49].
The detected keypoints are used to sample their corresponding descriptors $d\in\mathbb{R}^{256}$ from $D$. In Sec. 3.1, we provide an overview of DKD [49] and deformable attention [52]. In Sec. 3.2, we introduce the network architecture. In Sec. 3.3, we discuss how sparse and dense matching are obtained. Finally, the training details and loss functions are presented in Sec. 3.4.
@@ 3.1.1 Differentiable Keypoint Detection
where $t_{det}$ is the temperature. Then $s^{\prime}(i,j)$ represents the probability of $[i,j]^{T}$ to be a keypoint within the local window. Thus, the expected position of the keypoint in the local window can be given by integral regression [43, 14]:
$[\hat{i},\hat{j}]_{soft}^{T}=\sum_{0\leq i,j<N}s^{\prime}(i,j)[i,j]^{T}.$ (2)
@@ 3.1.2 Deformable Attention
$\mathbf{x}_{d}(\mathbf{z}_{q},\mathbf{p}_{q},\mathbf{x})=\sum_{m=1}^{M}\mathbf{W}_{m}A_{mq},\vspace{-1 em}$ (4)
@@
$A_{mq}=\sum_{k=1}^{K}A_{mqk}\cdot\mathbf{W}^{\prime}_{m}\mathbf{x}(\mathbf{p}_{q}+\Delta\mathbf{p}_{mqk}),\vspace{-0.5 em}$ (5)
@@
Most of recent keypoint detection and description frameworks [10, 33, 49, 50] benefit from the use of multi-scale feature maps to capture fine details and broader spatial context. Deformable attention naturally extends to multi-scale feature maps [52], enabling the network to effectively adapt to features at different scales. Similar to Eq. 4, multi-scale deformable attention feature $\mathbf{x}_{msd}$ can be calculated similar to Eq. 4 and Eq. 7
$\mathbf{x}_{msd}(\mathbf{z}_{q},\mathbf{p}_{q},\{\mathbf{x}^{l}\}_{l=1}^{L})=\sum_{m=1}^{M}\mathbf{W}_{m}A^{\prime}_{mq},\vspace{-1em}$ (6)
@@ 3.2 Network Architecture
Using a ResNet-50 [15] as the CNN backbone, $\mathcal{F}_{res}$ extracts feature maps at 4 scales, each with resolutions of $1/4,1/8,1/16,1/32$ of the original image. An additional feature map at scale of $1/64$ is then included by applying a CNN on the last feature map. These features are positionally encoded and fed into $\mathcal{F}_{e}$ with multi-scale deformable attention [52], enabling the network to dynamically attend to relevant spatial locations across multiple scales. We use 4 encoder layers each with 8 attention heads and each head samples 8 points. The entire process of the Descriptor Branch is illustrated in Fig. 2.
@@
The keypoint branch is dedicated solely to detecting accurate keypoints without being influenced by descriptor estimation. $\mathcal{F}_{cnn}$ extracts feature maps at 4 different resolutions, each with a resolution of $1/1,1/2,1/8,$ and $1/32$ of the original image and feature dimension of 32. These features are subsequently upsampled to match the original image resolution and concatenated to produce a feature map of size $H\times W\times 128$. Then $\mathcal{F}_{s}$ is used to estimate the scoremap. DKD [49] is applied on the scoremap to perform non-maximum suppression and detect the sub-pixel-accurate keypoints.  Fig. 2 depicts the entire process of the Keypoint Branch.

### v5

@@ 1 Introduction
Keypoint detection and description are central to numerous 3D computer vision tasks, including structure-from-motion (SfM), visual localization, and simultaneous localization and mapping (SLAM). These tasks depend on reliable keypoints and descriptors to establish accurate tracks, which are crucial for downstream algorithms to interpret spatial relationships—a requirement in fields from robotics to augmented reality [kerbl3Dgaussians]. When applied to real-world tasks, these systems face challenges such as significant viewpoint shifts, lighting variations, and scale changes.
@@
Historically, feature detection and description methods relied on rule-based techniques such as SIFT , SURF, and ORB [Lowe:2004:DIF:993451.996342, SURF2008, ORB2011]. However, these methods oftenf struggled with the aforementioned challenges. To better address these issues, recent works [superpoint2018, gleize2023silksimplelearned, tyszkiewicz2020disk, Zhao2022ALIKE] leverage deep neural networks based on data-driven approaches which often extract more robust and discriminative descriptors than the rule-based ones. Despite the enhanced robustness that learning-based methods have, they often rely on regular convolutions to encode images [superpoint2018, tyszkiewicz2020disk, gleize2023silksimplelearned, potje2024cvpr, Zhao2022ALIKE], which ignore the geometric invariance and long-distance awareness critical for robust description under challenging conditions. Subsequently, [luo2020aslfeat, LIFT2016, liu2019giftlearningtransformationinvariantdense] have partially addressed geometric invariance by estimating the scale and orientation of keypoint descriptors to model affine transformations. ALIKED [zhao2023aliked] and ASLFeat [luo2020aslfeat] advanced this approach by using deformable convolutions capable of modeling any geometric transformations. However, these methods are limited to modeling transformations within local windows; features learned with the localized kernels of convolution operators fall short of learning important visual cues that depend on long-range relationships (e.g. vanishing lines).
@@
In this paper, we focus on feature detection and description in challenging scenarios that haven’t been effectively addressed in prior works – specifically, extracting reliable and discriminative keypoints and descriptors across images with large camera baselines, significant illumination variations, and scale changes. We therefore propose RDD: a novel two-branch architecture for keypoint detection and description that can model both geometric invariance and global context at the same time. In particular, we design two dedicated network branches: a fully convolutional architecture to perform keypoint detection and a transformer-based architecture for extracting descriptors: the objectives of keypoint detection and description have previously been found not to align perfectly [li2022decoupling]. We observe that the convolutional neural network excels in detecting sub-pixel keypoints thanks to its expressiveness, whereas a transformer-based architecture is particularly suited for learning features that captures global context and possess geometric invariance through its self-attention mechanism [vaswani2017attention, dosovitskiy2021imageworth16x16words]. However, as self-attention layers attend to all spatial locations in the image, they significantly increase computational overhead and may reduce descriptor discriminability. To address this, we have adopted the deformable attention [zhu2020deformable] for keypoint description, allowing our designed network to selectively focus on key locations, which greatly reduces time complexity while maintaining the capability to learn geometric invariance and global context.
@@
The two-branch architecture of RDD allows specialized architectures to learn keypoint detection and description independently. Our experiments have found that such design choice accelerates convergence and leads to better overall performance, as proven in Sec. 4.4. Overall, RDD is designed to be robust, achieving competitive performance on challenging imagery and also capable of semi-dense matching Fig. 1. This enhanced capability improves accuracy and stability in 3D vision applications, such as structure-from-motion (SfM) [chen2024geometryawarefeaturematchinglargescale] and relative camera pose estimation under difficult imaging conditions. Our approach outperforms current state-of-the-art methods for keypoint detection and description on standard benchmarks, including MegaDepth-1500 [dusmanu2019d2net], HPatches [hpatches_2017_cvpr], and Aachen-Day-Night [sattler2018benchmarking6dofoutdoorvisual], across various tasks. However, to the best of our knowledge, no existing benchmarks accurately capture the aftermentioned challenging scenarios. To further validate the robustness of our method and provide a more comprehensive evaluation, we have collected and derived two additional benchmark datasets. The main contributions of our work are:
@@
  • Dataset contribution: MegaDepth-View, derived from MegaDepth [li2018megadepth], specifically focuses on challenging image pairs with large viewpoint and scale variations. Air-to-Ground dataset provides diverse aerial and ground cross view imagery for training and evaluation, enhancing robustness and performance and providing a new evaluation setting.
Figure 2: An overview of our network architecture. Descriptor Branch $\mathcal{F}_{D}$ and Keypoint Branch $\mathcal{F}_{K}$ process an input image $I\in\mathbb{R}^{H\times W\times 3}$ independently. Descriptor Branch: 4 layers of multiscale feature maps $\{\mathbf{x}_{res}^{l}\}_{l=1}^{L}$ are extracted by passing II through ResNet-50 [he2015deepresiduallearningimage] $\mathcal{F}_{res}$. Additional feature map at scale of $1/64$ is added by applying a simple CNN on the last feature map and then they are feed to a transformer encoder $\mathcal{F}_{e}$ with positional embeddings [vaswani2017attention]. We up-sample all feature maps output by $\mathcal{F}_{e}$ to size $H/k\times W/k$ where $k=4$ is the patch size. Feature maps are then summed together to generate the dense descriptor map $D$. A classification head $\mathcal{F}_{m}$ is applied to the descriptor map to estimate a matchability map $M$. Keypoint Branch: II passes through a lightweight CNN with residual connection [he2015deepresiduallearningimage] $\mathcal{F}_{cnn}$ to capture multi-scale features $\{\mathbf{x}_{cnn}^{l}\}_{l=1}^{L}$. Features are then up-sampled to size $H\times W$ and concatenated to generate a feature map of $H\times W\times 64$. A score map $S$ is estimated by a classification head $\mathcal{F}_{s}$. Final sub-pixel keypoints are detected using DKD [Zhao2022ALIKE].
@@ 2.1 Geometric Invariant Descriptors
In geometric invariance modeling, previous methods have primarily focused on two key aspects: scale and orientation. In traditional hand-crafted methods, SIFT [siftflow] estimates a keypoint’s scale and orientation by analyzing histograms of image gradients and subsequently extracts and normalizes image patches around the keypoint to construct scale- and orientation-invariant descriptors. Similarly, ORB [ORB2011] computes orientation based on the keypoint’s center of mass, followed by rotation of the image patches to ensure orientation invariance.
Learning-based methods approach geometric invariance through two primary strategies. The majority of these methods [superpoint2018, tyszkiewicz2020disk, dusmanu2019d2net, r2d2, gleize2023silksimplelearned, potje2024cvpr] relies heavily on data augmentation techniques to achieve scale and orientation invariance. On the other hand, some learning-based approaches explicitly model scale and orientation. For instance, LIFT [LIFT2016] mimics SIFT [siftflow] by detecting keypoints, estimating their orientation, and extracting descriptors with different neural networks. Other methods, such as AffNet [mishkin2018repeatabilityenoughlearningaffine] and LF-Net [ono2018lfnetlearninglocalfeatures], predict affine transformation parameters and apply these transformations via Spatial Transformer Networks [jaderberg2016spatialtransformernetworks] to produce affine-invariant descriptors. These methods generally assume a predefined geometric transformation, typically affine transformations. More recent advancements, such as ASLFeat [luo2020aslfeat] and ALIKED [zhao2023aliked], employ deformable convolutions to directly extract geometric-invariant features, offering a more flexible and adaptive approach to handling geometric variations.
@@ 2.2 Decoupling Keypoint and Descriptor Learning
In traditional handcrafted methods for feature detection and description, such as SIFT [siftflow] and ORB [ORB2011], the detection and description processes are performed independently. This decoupled approach was also common in early learning-based methods [barrosolaguna2019keynetkeypointdetectionhandcrafted, mishchuk2018workinghardknowneighbors], where keypoint detection and description were treated as separate tasks. However, later works [dusmanu2019d2net, superpoint2018, gleize2023silksimplelearned, r2d2] proposed an end-to-end framework, jointly optimizing keypoint detection and descriptor extraction. This joint learning approach ensures that the detected keypoints and their descriptors complement each other. Despite the advantages,  [li2022decoupling] highlighted a potential drawback of joint training, observing that poorly learned descriptors could negatively impact keypoint detection quality. Subsequently, DeDoDe[edstedt2024dedode] demonstrated that decoupling keypoint detection and description into two independent networks can significantly improve performance. Similarly, XFeat [potje2024cvpr] showed that using task-specific networks provides a better balance between efficiency and accuracy. Inspired by these recent insights, RDD consists of two separate branches for independent keypoint detection and description. We also formulate a training strategy that optimize the descriptor branch first and then the keypoint branch. This design enables both sparse and dense matching and ensures optimized performance across various scenarios.
@@ 2.3 Efficient Attention Mechanisms
The quadratic complexity of multi-head attention [vaswani2017attention] leads to long training schedules and limited spatial resolution for tractability. For 2D images, LR-Net [hu2019local] proposed a local relation layer, computing attention in a fixed local window. Linear Transformer [katharopoulos2020transformersrnnsfastautoregressive] proposes to reduce the computation complexity of vanilla attention to linear complexity by substituting the exponential kernel with an alternative kernel function. Deformable attention [zhu2020deformable] further allows the attention window to depend on the query features. With a linear projection predicting a deformable set of displacements from the center and their weight, each pixel in the feature map may attend to a fixed number of pixels of arbitrary distance. Deformable attention [zhu2020deformable] is well-suited for our objectives, enabling efficient learning of global context, making it a fitting choice for our approach.
@@ 3 Methods
Our method extracts descriptors and keypoints separately using a descriptor branch and a keypoint branch, as shown in Fig. 2. For an input image $I\in\mathbb{R}^{H\times W\times 3}$, the descriptor branch first extracts multi-scale feature maps $\{x_{res}^{l}\}_{l=1}^{L}$ from $I$ using ResNet-50 [he2015deepresiduallearningimage], denoted as $\mathcal{F}_{res}$, and then estimates a descriptor map $D\in\mathbb{R}^{H/k\times W/k}$ through a deformable transformer encoder [zhu2020deformable], denoted as $\mathcal{F}_{e}$, where $k=4$ is the patch size. The matchability map $M\in\mathbb{R}^{H/k\times W/k}$ is then estimated from $D$ with a classification head $\mathcal{F}_{m}$. The keypoint branch $\mathcal{F}_{K}$ estimates a scoremap $S\in\mathbb{R}^{H\times W}$ by passing image $I$ through a lightweight CNN with residual connection [he2015deepresiduallearningimage] $\mathcal{F}_{cnn}$ and a classification head $\mathcal{F}_{s}$, then sub-pixel keypoints $p_{s}=[x,y]^{T}$ are detected with differentiable keypoint detection (DKD) [Zhao2022ALIKE].
The detected keypoints are used to sample their corresponding descriptors $d\in\mathbb{R}^{256}$ from $D$. In Sec. 3.1, we provide an overview of DKD [Zhao2022ALIKE] and deformable attention [zhu2020deformable]. In Sec. 3.2, we introduce the network architecture. In Sec. 3.3, we discuss how sparse and dense matching are obtained. Finally, the training details and loss functions are presented in Sec. 3.4.
@@ 3.1.1 Differentiable Keypoint Detection
where $t_{det}$ is the temperature. Then $s^{\prime}(i,j)$ represents the probability of $i,j]^{T}$ to be a keypoint within the local window. Thus, the expected position of the keypoint in the local window can be given by integral regression [sun2018integralhumanposeregression, gu2021removing]:
$\hat{i},\hat{j}]_{soft}^{T}=\sum_{0\leq i,j<N}s^{\prime}(i,j)[i,j]^{T}.$ (2)
@@ 3.1.2 Deformable Attention
$\mathbf{x}_{d}(\mathbf{z}_{q},\mathbf{p}_{q},\mathbf{x})=\sum_{m=1}^{M}\mathbf{W}_{m}A_{mq},\vskip-10.00002pt$ (4)
@@
$A_{mq}=\sum_{k=1}^{K}A_{mqk}\cdot\mathbf{W}^{\prime}_{m}\mathbf{x}(\mathbf{p}_{q}+\Delta\mathbf{p}_{mqk}),\vskip-5.0pt$ (5)
@@
Most of recent keypoint detection and description frameworks [edstedt2024dedode, potje2024cvpr, Zhao2022ALIKE, zhao2023aliked] benefit from the use of multi-scale feature maps to capture fine details and broader spatial context. Deformable attention naturally extends to multi-scale feature maps [zhu2020deformable], enabling the network to effectively adapt to features at different scales. Similar to Eq. 4, multi-scale deformable attention feature $\mathbf{x}_{msd}$ can be calculated similar to Eq. 4 and Eq. 7
$\mathbf{x}_{msd}(\mathbf{z}_{q},\mathbf{p}_{q},\{\mathbf{x}^{l}\}_{l=1}^{L})=\sum_{m=1}^{M}\mathbf{W}_{m}A^{\prime}_{mq},\vskip-10.00002pt$ (6)
@@ 3.2 Network Architecture
Using a ResNet-50 [he2015deepresiduallearningimage] as the CNN backbone, $\mathcal{F}_{res}$ extracts feature maps at 4 scales, each with resolutions of $1/4,1/8,1/16,1/32$ of the original image. An additional feature map at scale of $1/64$ is then included by applying a CNN on the last feature map. These features are positionally encoded and fed into $\mathcal{F}_{e}$ with multi-scale deformable attention [zhu2020deformable], enabling the network to dynamically attend to relevant spatial locations across multiple scales. We use 4 encoder layers each with 8 attention heads and each head samples 8 points. The entire process of the Descriptor Branch is illustrated in Fig. 2.
@@
The keypoint branch is dedicated solely to detecting accurate keypoints without being influenced by descriptor estimation. $\mathcal{F}_{cnn}$ extracts feature maps at 4 different resolutions, each with a resolution of $1/1,1/2,1/8,$ and $1/32$ of the original image and feature dimension of 32. These features are subsequently upsampled to match the original image resolution and concatenated to produce a feature map of size $H\times W\times 128$. Then $\mathcal{F}_{s}$ is used to estimate the scoremap. DKD [Zhao2022ALIKE] is applied on the scoremap to perform non-maximum suppression and detect the sub-pixel-accurate keypoints.  Fig. 2 depicts the entire process of the Keypoint Branch.